OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? Answer 4x + y = 1 x - 4y = -13 4x - y = 1 4x - y = -7 im confused i tried and got D 4x - y = -7

OpenStudy (anonymous):

(y2-y1)/(x2-x1) to get the slope

OpenStudy (anonymous):

then equation of the line y-y1=m(x-x1) m is the slope

sam (.sam.):

m=4 (1, 11) y-y1=m(x-x1) y=4x+7

OpenStudy (anonymous):

y=4x+7

OpenStudy (anonymous):

i guess this the answer u want 4x - y = -7 , just rearrange it to be y=4x+7

OpenStudy (anonymous):

that's how equation of a line is represented